Evaluate the integral by reversing the order of integration. 6. S. Brywą dy de 3.xy3/2 dy de
Answer #1

y y=x У= 4/ х=0 (1,1). C = 1 2.. I from the graph we can change the order of the integration easily I J 3 xyh dy du = 3х у y=0 x=y4 1 hun Y 3/2 JU dze dy a